A student wants to prepare 250 mL of a 0.035 mol L-1 solution from the primary standard, anhydrous sodium carbonate. Calculate the mass of anhydrous sodium carbonate that is needed.
1
Expert's answer
2019-04-26T00:57:32-0400

n(Na2CO3) = C(Na2CO3) × V(Na2CO3) = 0.035 mol L-1 × 0.25 L = 0.00875 mol.

m(Na2CO3) = n(Na2CO3) × M(Na2CO3) = 0.00875 mol × (2×23 + 12 + 3×16) = 0.9275 g = 927.5 mg.